# OpenFACADES: An Open Framework for Architectural Caption and Attribute Data Enrichment via Street View Imagery

### v1

@@ Abstract
Building properties, such as height, usage, and material composition, play a crucial role in spatial data infrastructures, supporting applications such as energy simulation, risk assessment, and environmental modeling. Despite their importance, comprehensive and high-quality building attribute data remain scarce in many urban areas. Recent advances have enabled the extraction and tagging of objective building attributes using remote sensing and street-level imagery. However, establishing a method and pipeline that integrates diverse open datasets, acquires holistic building imagery at scale, and infers comprehensive building attributes remains a significant challenge. Among the first, this study bridges the gaps by introducing OpenFACADES, an open framework that leverages multimodal crowdsourced data to enrich building profiles with both objective attributes and semantic descriptors through multimodal large language models. Our methodology proceeds in three major steps. First, we integrate street‐level image metadata from Mapillary with OpenStreetMap geometries via isovist analysis, effectively identifying images that provide suitable vantage points for observing target buildings. Second, we automate the detection of building facades in panoramic imagery and tailor a reprojection approach to convert objects into holistic perspective views that approximate real-world observation. Third, we introduce an innovative approach that harnesses and systematically investigates the capabilities of open-source large vision-language models (VLMs) for multi-attribute prediction and open-vocabulary captioning in building-level analytics, leveraging a globally sourced dataset of 30,180 labeled images from seven cities. Evaluation shows that fine-tuned VLM excel in multi-attribute inference, outperforming single-attribute computer vision models and zero-shot ChatGPT-4o. Further experiments confirm its superior generalization and robustness across culturally distinct regions and varying image conditions. Finally, the model is applied for large-scale building annotation, generating a dataset of 1.2 million images for half a million buildings. This open‐source framework enhances the scope, adaptability, and granularity of building‐level assessments, enabling more fine‐grained and interpretable insights into the built environment. Our dataset and code are available openly at: https://github.com/seshing/OpenFACADES.
@@ 1 Introduction
Buildings, as prominent artifacts within urban settings, serve as vital indicators of the management, transformation, and overall dynamism of the built environment. Their physical characteristics — including geometry, height, function, material, condition, and style — are the key parameters that not only support sustainable urban development but also reflect economic progress and cultural evolution over time (Biljecki et al., 2021). Such rich building-level data has been instrumental in a range of applications, such as urban climate simulations for improved environmental planning (Creutzig et al., 2019), building energy modeling for resource optimization (Kumar et al., 2018; Roth et al., 2020), estimation of urban material stocks for the circular economy (Raghu et al., 2023), and disaster impact assessments to inform effective response and recovery efforts (Westrope et al., 2014). Moreover, these data support more nuanced analyses of population distributions (Schug et al., 2021), socio-economic conditions (Feldmeyer et al., 2020), as well as deeper understanding of the impact on human behaviors (Wang et al., 2016) and public perception (Liang et al., 2024). Hence, more comprehensive and openly accessible geospatial data on building information can enable the formulation of nuanced urban planning policies, fostering locally informed and globally connected approaches to efficiently support urban resilience and sustainability (Elmqvist et al., 2019).
@@
Traditionally, obtaining building attributes has involved expert evaluation, government records, or crowdsourced labeling, which often require field studies. This approach limits coverage and efficiency, leaving many buildings without detailed information. Although platforms like OpenStreetMap (OSM) and government databases now contain diverse urban information, the incompleteness and uneven geographical distribution of global building features hinder their usability across larger regions (Biljecki et al., 2023; Lei et al., 2023; Herfort et al., 2023). With their rapid development, remote sensing-based methods have become a standard approach for extracting building information from aerial and satellite imagery, including attributes such as building height (Wu et al., 2023b; Frantz et al., 2021), and types (Du et al., 2015; Zhao et al., 2019). Remote sensing technologies offer significant advantages, such as large-scale data coverage, reduced dependency on ground surveys, and the ability to monitor urban changes over time with high spatial and temporal resolution. In parallel, machine learning methods that leverage geometric and built environment information have been widely applied to enhance the coverage and accuracy of building data (Roy et al., 2023; Milojevic-Dupont et al., 2023; Lei et al., 2024; Wang et al., 2024c). Despite the advancements, the top-down observation of these technologies poses inherent challenges for object-based building evaluation, as the vertical dimension of buildings holds critical and detailed information that is often difficult to capture.
@@
The emergence of easily accessible Street View Imagery (SVI) has transformed the way buildings are analyzed, providing a ground-level, bottom-up perspective that captures architectural details often obscured in aerial or satellite imagery (Biljecki and Ito, 2021; Gaw et al., 2022; Zhang et al., 2024). Leveraging this capability, numerous studies have integrated deep learning with SVI to extract and profile various building attributes, including height (Yan and Huang, 2022; Fan et al., 2024), type and usage (Kang et al., 2018; Zhao et al., 2021; Ramalingam and Kumar, 2023), architectural style (Lindenthal and Johnson, 2021; Sun et al., 2022b), and facade materials (Xu et al., 2023; Raghu et al., 2023; Chen et al., 2024a). Beyond building profiling, these integrations also support a range of practical applications, including risk assessment (Pelizari et al., 2021; Wang et al., 2021), refinement of 3D building models (Zhang et al., 2021), and building energy efficiency estimation (Sun et al., 2022a; Mayer et al., 2023). These advancements have significantly contributed to SVI-based urban studies, enabling fine-grained, large-scale geospatial analyses.
@@
Despite advancements in SVI-based methods for inferring building attributes, various challenges limit scalability and adaptability: (1) existing datasets struggle with uncertainty due to limited angular coverage in perspective views or distortions in panoramic images, hindering comprehensive observations; (2) reliance on proprietary data restricts accessibility, transparency, and adaptability, with ambiguous licensing further limiting research utility and inclusivity (Helbich et al., 2024); (3) while some efforts align visual data with geolocation, annotations often focus on isolated attributes, requiring separate models. Multi-task learning has been explored (Chen et al., 2022), but class diversity and scalability remain constrained. Rigid annotation schemes further prevent adaptation to emergent characteristics like mixed-use functions or hybrid materials, limiting the ability to capture architectural complexity for more inclusive and interpretable analyses. As a consequence of these challenges, the volume of building datasets derived from SVI that offer a holistic view of structures, fully rely on open datasets, and support comprehensive architectural insights remains quite limited.
@@
Vision-language models (VLMs) — which integrate advanced computer vision (CV) and natural language processing — have demonstrated the ability to interpret complex visual relationships, reason about scenes, and generate coherent, semantically rich descriptions (Li et al., 2024a). In the remote sensing domain, vision-language tasks have demonstrated promise for multi-scale feature understanding, multi-task learning, and applications such as visual question answering, image captioning, and semantic segmentation (Zia et al., 2022; Hu et al., 2023; Dong et al., 2024; Wang et al., 2024a). More recently, multimodal large language models (MLLMs) have advanced these capabilities by integrating deep contextual and semantic representations learned from massive, multimodal datasets, thereby enabling more nuanced and precise interpretations of visual data. This versatility highlights their potential to serve as foundational instruments in SVI-based building research, by enhancing the characterization of building properties, streamlining multi-task learning, and transcending predefined label sets in the analysis of facade features.
@@
To advance fine-grained, bottom-up observations of buildings, we propose an open framework, OpenFACADES, that enriches a variety of building properties from a street-level perspective by leveraging multimodal crowdsourced inputs and open-source MLLMs. First, we utilize fully open-source building footprints and SVI to perform visibility simulations that geospatially align and integrate visible building geometries with corresponding SVI shooting locations. Second, we introduce an innovative pipeline that detects individual buildings based on their visible angles and acquires holistic building images using a custom image reprojection method. Third, we assemble one of the largest global, multi-attribute building image datasets by combining crowdsourced building attributes with high-quality text descriptions generated by state-of-the-art MLLMs. Leveraging this dataset, we are among the first to introduce tailored MLLMs for building profiling through multi-task learning, encompassing both single- and multi-attribute prediction tasks as well as open-vocabulary captioning. Furthermore, we present an in-depth comparative analysis of model performance across various hyperparameter settings, cross-city generalization scenarios, and image quality variations, discussing key challenges identified in prior urban studies (Sun et al., 2022b; Hou et al., 2025; Hou and Biljecki, 2022). These experiments serve as the first systematic demonstration of applying multimodal large language models (MLLMs) in building-level studies, paving the way for their broader adoption in urban research.
@@
  • Compiled an open global building dataset, consisting of (1) 30,180 individual building images from seven cities across three continents, annotated with attribute labels from OSM and text descriptions generated by ChatGPT-4o; and (2) large-scale automated annotations on 1.2 million images covering over half a million buildings. Each image is linked to its geospatial location and enriched with diverse attributes (e.g., building type, number of floors, age, and surface material) along with detailed textual descriptions. This forms the OpenFACADE dataset, one of the largest such resources, spanning multiple urban morphologies.
@@ 2.1 Existing street-level building datasets
With advances in geospatial artificial intelligence technologies, research in recent years has increasingly leveraged remote sensing datasets such as high-resolution satellite and aerial imagery, and LiDAR to enhance urban development and planning applications. These datasets enable object-based image analysis, pixel-based classification, and semantic segmentation of urban structures, providing critical insights for land use mapping, urban morphology analysis, and spatiotemporal change detection. As key urban components, buildings have spurred the creation of domain-specific datasets and methodologies to support applications such as urban sustainability evaluation through rooftop attributes extraction (Wu and Biljecki, 2021), infrastructure management via automated land cover classification (Boguszewski et al., 2021), and disaster management through assessing damage (Gupta et al., 2019; Li et al., 2025a).
@@
SVI, rapidly emerging as a prominent proximal remote sensing data source, has been leveraged to generate spatially enriched urban datasets that facilitate fine-grained semantic understanding of complex urban scenes (Biljecki and Ito, 2021). Among these, building-centric SVI datasets enable facade-level feature extraction, offering images that capture textural, material, and architectural features of building exteriors for environmental modeling. Building age and architectural style — key indicators of urban morphological evolution — have long been studied for their correlations with building thermal performance (Tooke et al., 2014; Aksoezen et al., 2015; Nouvel et al., 2017) and influence on pricing models in real estate (Zietz et al., 2008; Lindenthal and Johnson, 2021). Recent advances include the work of Sun et al. (2022b), which applies deep convolutional neural networks (CNNs) to classify buildings in Amsterdam, the Netherlands, into architectural periodization categories (e.g., revival, postwar). Material characterization (Xu et al., 2023; Chen et al., 2024a), another aspect critical for building energy simulation (Nouvel et al., 2017), also supports circular economy objectives by enabling lifecycle material tracking (Raghu et al., 2023) and risk assessment (Wang et al., 2021). Among these efforts, Raghu et al. (2023) employ a multi-city material categories (brick, stucco, etc.) using geotagged SVI perspective views, aligning visual patterns with ground-truth material information for scalable building classification. Combining the aspects of building age and material, Ogawa et al. (2023) introduced a method to detect and geolocate buildings from panoramic images, automatically annotating them with objective building data in Kobe, Japan.
@@
Furthermore, building type or usage — a critical attribute in urban remote sensing and land use classification frameworks — is another important aspect in street-level research (Kang et al., 2018; Zhao et al., 2021; Lindenthal and Johnson, 2021; Ramalingam and Kumar, 2023; Li et al., 2025b). A seminal work by Kang et al. (2018) introduces the BIC_GSV dataset, a multi-city geospatial database of 19,658 SVI-derived building facades categorized into eight classes (e.g., apartment, church, garage, etc.) across North America. These ground truth labels are generated through view-direction-aligned spatial joins with OSM building footprints, enabling parcel-scale urban pattern analysis. Advancing this, Zhao et al. (2021) developed the BEAUTY dataset, which extends BIC_GSV by incorporating both SVI-based land use classification (e.g., residential, commercial, etc.) and multi-class building detection. Other similar research frameworks have also been applied to large-scale urban studies, integrating additional building attributes such as floor number estimation, abandoned house detection, and seismic risk assessment (Iannelli and Dell’Acqua, 2017; Zou and Wang, 2021; Rosenfelder et al., 2021; Pelizari et al., 2021; Ghione et al., 2022). These workflows not only enable location-based building retrieval but also demonstrate cross-modal alignment of SVI with open geospatial building footprints, optimizing multi-source training data generation for CV pipelines.
@@
However, several challenges still remain in street-level building research, limiting the scalability and adaptability of current approaches. First, although many efforts have aligned visual information with building geolocation (Kang et al., 2018; Sun et al., 2022c; Ogawa et al., 2023), they are often either reliant on perspective views with restricted angular coverage, limiting visibility of upper building elements, or on panoramic images prone to severe distortions, misaligning with actual observations. Second, while various SVI-based building datasets have been established, their dependence on data derived from proprietary platforms introduces limitations related to accessibility, transparency, and adaptability. The ambiguous licensing terms of such datasets further constrain their utility for diverse research applications and compromise the integrity of work built upon them, thereby hindering inclusivity within the research community (Helbich et al., 2024). In a recent trend, crowdsourced SVI platforms have garnered attention in urban studies by producing diverse, publicly accessible imagery, prompting efforts to expand dataset coverage and customized applications. Examples include annotating points of interest (Zarbakhsh and McArdle, 2023), image status (Hou et al., 2024), human perception (Yang et al., 2025), and road surface type (Kapp et al., 2025). Among these, Hou et al. (2024) curate a manually labeled dataset to assess 10 million crowdsourced SVIs from 688 cities, enriched with metadata such as platform, weather, and lighting conditions, while Kapp et al. (2025) utilize OSM tags and ChatGPT-4o to label and amplify underrepresented road surface classes, resulting in 9,122 labeled images. These initiatives illustrate the potential of crowdsourced data for broad, inclusive urban analyses.

### v2

@@ Abstract
Building properties, such as height, usage, and material, play a crucial role in spatial data infrastructures, supporting various urban applications. Despite their importance, comprehensive building attribute data remain scarce in many urban areas. Recent advances have enabled the extraction of objective building attributes using remote sensing and street-level imagery. However, establishing a pipeline that integrates diverse open datasets, acquires holistic building imagery, and infers comprehensive building attributes at scale remains a significant challenge. Among the first, this study bridges the gaps by introducing OpenFACADES, an open framework that leverages multimodal crowdsourced data to enrich building profiles with both objective attributes and semantic descriptors through multimodal large language models. First, we integrate street-level image metadata from Mapillary with OpenStreetMap geometries via isovist analysis, identifying images that provide suitable vantage points for observing target buildings. Second, we automate the detection of building facades in panoramic imagery and tailor a reprojection approach to convert objects into holistic perspective views that approximate real-world observation. Third, we introduce an innovative approach that harnesses and investigates the capabilities of open-source large vision-language models (VLMs) for multi-attribute prediction and open-vocabulary captioning in building-level analytics, leveraging a globally sourced dataset of 31,180 labeled images from seven cities. Evaluation shows that fine-tuned VLM excel in multi-attribute inference, outperforming single-attribute computer vision models and zero-shot ChatGPT-4o. Further experiments confirm its superior generalization and robustness across culturally distinct region and varying image conditions. Finally, the model is applied for large-scale building annotation, generating a dataset of 1.2 million images for half a million buildings. This open‐source framework enhances the scope, adaptability, and granularity of building‐level assessments, enabling more fine‐grained and interpretable insights into the built environment. Our dataset and code are available openly at: https://github.com/seshing/OpenFACADES.
@@ 1 Introduction
Buildings, as prominent artifacts within urban settings, serve as vital indicators of the management, transformation, and overall dynamism of the built environment. Their physical characteristics, including geometry, height, function, material, condition, and style, are the key parameters that not only support sustainable urban development but also reflect economic progress and cultural evolution over time (Biljecki et al., 2021). Such rich building-level data has been instrumental in a range of applications, such as urban climate simulations for improved environmental planning (Creutzig et al., 2019), building energy modeling for resource optimization (Kumar et al., 2018; Roth et al., 2020), estimation of urban material stocks for the circular economy (Raghu et al., 2023), and disaster impact assessments to inform effective response and recovery efforts (Westrope et al., 2014). Moreover, these data support more nuanced analyses of population distributions (Schug et al., 2021), socio-economic conditions (Feldmeyer et al., 2020), as well as deeper understanding of the impact on human behaviors (Wang et al., 2016) and public perception (Liang et al., 2024). Hence, more comprehensive and openly accessible geospatial data on building can enable the formulation of nuanced urban planning policies, fostering locally informed and globally connected approaches to efficiently support urban resilience and sustainability (Elmqvist et al., 2019).
@@
Traditionally, obtaining building attributes has involved expert evaluation, government records, or crowdsourced labeling, which often require field studies. This approach limits coverage and efficiency, leaving many buildings without detailed information. Although platforms like OpenStreetMap (OSM) and government databases now contain diverse urban information, the incompleteness and uneven geographical distribution of global building features hinder their usability across larger regions (Biljecki et al., 2023; Milojevic-Dupont et al., 2023; Lei et al., 2023; Herfort et al., 2023; Florio et al., 2025). With their rapid development, remote sensing-based methods have become a standard approach for extracting building information from aerial and satellite imagery, including attributes such as building height (Wu et al., 2023b; Frantz et al., 2021), and types (Du et al., 2015; Zhao et al., 2019). Remote sensing provides broad coverage, reduces reliance on ground surveys, and enables high-resolution tracking of urban changes over time. In parallel, machine learning methods that leverage geometric and built environment information have been widely applied to enhance the coverage and accuracy of building data (Roy et al., 2023; Nachtigall et al., 2023; Lei et al., 2024; Wang et al., 2024c). Despite the advancements, the top-down perspective poses inherent challenges, as critical vertical details of structures remain difficult to capture from overhead imagery.
@@
The emergence of easily accessible Street View Imagery (SVI) has transformed the way buildings are analyzed, providing a ground-level, bottom-up perspective that captures architectural details often obscured in aerial or satellite imagery (Biljecki and Ito, 2021; Gaw et al., 2022; Zhang et al., 2024a). Leveraging this capability, numerous studies have integrated deep learning with SVI to extract and profile various building attributes, including height (Yan and Huang, 2022; Fan et al., 2024), type and usage (Kang et al., 2018; Zhao et al., 2021; Ramalingam and Kumar, 2023), architectural style (Lindenthal and Johnson, 2021; Sun et al., 2022b), and facade materials (Xu et al., 2023; Raghu et al., 2023; Chen et al., 2024a). Beyond building profiling, these integrations also support a range of practical applications, including risk assessment (Pelizari et al., 2021; Wang et al., 2021), refinement of 3D building models (Zhang et al., 2021), and building energy efficiency estimation (Sun et al., 2022a; Mayer et al., 2023). These advancements have significantly contributed to SVI-based urban studies, enabling fine-grained, large-scale geospatial analyses.
@@
Despite advancements in SVI-based methods for inferring building attributes, various challenges limit scalability and adaptability: (1) existing datasets struggle with uncertainty due to limited angular coverage in perspective views or distortions in panoramic images, hindering comprehensive observations; (2) reliance on proprietary data restricts accessibility, transparency, and adaptability, with ambiguous licensing further limiting research utility and inclusivity (Helbich et al., 2024); (3) while some efforts align visual data with geolocation, annotations often focus on isolated attributes, requiring separate models. Multi-task learning has been explored (Chen et al., 2022), but class diversity remain constrained, limiting the ability to capture architectural complexity for more inclusive and interpretable analyses. Consequently, SVI-based building datasets offering holistic structural perspectives, fully open data, and comprehensive architectural insights remain scarce.
@@
Vision-language models (VLMs), uniting computer vision (CV) and natural language processing, have demonstrated the ability to interpret complex visual relationships, reason about scenes, and generate semantically rich descriptions (Li et al., 2024a). In the remote sensing domain, vision-language tasks have demonstrated promise for multi-scale feature understanding, multi-task learning, and applications such as visual question answering, image captioning, and semantic segmentation (Zia et al., 2022; Hu et al., 2025; Dong et al., 2024; Wang et al., 2024a). More recently, multimodal large language models (MLLMs) have advanced these capabilities by integrating deep contextual and semantic representations learned from massive, multimodal datasets, thereby enabling more nuanced and precise interpretations of visual data. This versatility highlights their potential to serve as foundational instruments in SVI-based building research, by enhancing the characterization of building properties, streamlining multi-task learning, and transcending predefined label sets in the analysis of facade features.
@@
To advance fine-grained, bottom-up observations of buildings, we propose an open framework, OpenFACADES, that enriches a variety of building properties from a street-level perspective by leveraging multimodal crowdsourced inputs and open-source MLLMs. First, we utilize open-source building footprints and SVI to perform visibility simulations that geospatially align building geometries with corresponding SVI shooting locations. Second, we introduce an innovative pipeline that detects individual buildings based on their visible angles and acquires holistic building images using a custom image reprojection method. Third, we assemble one of the largest global, multi-attribute building image datasets by combining crowdsourced building attributes with high-quality text descriptions generated by state-of-the-art MLLMs. Leveraging this dataset, we are among the first to introduce tailored MLLMs for building profiling through multi-task learning, encompassing both single- and multi-attribute prediction tasks as well as open-vocabulary captioning. Furthermore, we present an in-depth comparative analysis of model performance across various hyperparameter settings, cross-city generalization scenarios, and image quality variations.
@@
  • Compiled an open global building dataset, consisting of (1) 31,180 individual building images from seven cities across three continents, annotated with attribute labels from OSM and text descriptions generated by ChatGPT-4o; and (2) large-scale automated annotations on 1.2 million images covering over half a million buildings. Each image is linked to its geospatial location and enriched with diverse attributes (e.g., building type, number of floors, age, and surface material) along with detailed textual descriptions. This forms the OpenFACADES dataset, one of the largest such resources, spanning multiple urban morphologies.
@@ 2.1 Existing street-level building datasets
With advances in geospatial artificial intelligence technologies, research in recent years has increasingly leveraged remote sensing datasets such as high-resolution satellite and aerial imagery, and LiDAR to enhance urban applications. These datasets enable object-based image analysis, pixel-based classification, and semantic segmentation of urban structures, providing critical insights for land use mapping, urban morphology analysis, and spatiotemporal change detection. As key urban components, buildings have spurred the creation of domain-specific datasets and methodologies to support applications such as urban sustainability evaluation through rooftop attributes extraction (Wu and Biljecki, 2021), infrastructure management via automated land cover classification (Boguszewski et al., 2021), and disaster management through assessing damage (Gupta et al., 2019; Li et al., 2025a).
@@
SVI, rapidly emerging as a prominent proximal remote sensing data source, has been leveraged to generate spatially enriched urban datasets that facilitate fine-grained semantic understanding of complex urban scenes (Biljecki and Ito, 2021). Among these, building-centric SVI datasets enable facade-level feature extraction, offering images that capture textural, material, and architectural features of building exteriors for environmental modeling. For example, building age and architectural style have long been studied for their links to building thermal performance (Tooke et al., 2014; Aksoezen et al., 2015; Nouvel et al., 2017) and real estate pricing (Zietz et al., 2008; Lindenthal and Johnson, 2021). Recent advances include the work of Sun et al. (2022b), which applies deep convolutional neural networks (CNNs) to classify buildings in Amsterdam, the Netherlands, into architectural periodization categories (e.g., revival, postwar). Material characterization (Xu et al., 2023; Chen et al., 2024a), another aspect critical for building energy simulation (Nouvel et al., 2017), also supports circular economy objectives by enabling lifecycle material tracking (Raghu et al., 2023) and risk assessment (Wang et al., 2021). Among these efforts, Raghu et al. (2023) employ a multi-city material categories (brick, stucco, etc.) using geotagged SVI perspective views, aligning visual patterns with ground-truth material information for scalable building classification. Combining the aspects of building age and material, Ogawa et al. (2023) introduced a method to detect and geolocate buildings from panoramic images, automatically annotating them with objective building data in Kobe, Japan.
@@
Furthermore, building type or usage, a critical attribute in urban remote sensing and land use classification, is also central to street-level research (Kang et al., 2018; Zhao et al., 2021; Lindenthal and Johnson, 2021; Ramalingam and Kumar, 2023; Li et al., 2025b). A seminal work by Kang et al. (2018) introduces the BIC_GSV dataset, a multi-city geospatial database of 19,658 SVI-derived building facades categorized into eight classes (e.g., apartment, church, garage, etc.) across North America. These ground truth labels are generated through view-direction-aligned spatial joins with OSM building footprints, enabling parcel-scale urban pattern analysis. Advancing this, Zhao et al. (2021) developed the BEAUTY dataset, which extends BIC_GSV by incorporating both SVI-based land use classification (e.g., residential, commercial, etc.) and multi-class building detection. Other similar research frameworks have also been applied to large-scale urban studies, integrating additional building attributes such as floor number estimation, abandoned house detection, and seismic risk assessment (Iannelli and Dell’Acqua, 2017; Zou and Wang, 2021; Rosenfelder et al., 2021; Pelizari et al., 2021; Ghione et al., 2022). These workflows not only enable location-based building retrieval but also demonstrate cross-modal alignment of SVI with open geospatial building footprints.
@@
However, several challenges still remain in street-level building research, limiting the scalability and adaptability of current approaches. First, although many efforts have aligned visual information with building geolocation (Kang et al., 2018; Sun et al., 2022b; Ogawa et al., 2023), they are often either reliant on perspective views with restricted angular coverage, limiting visibility of upper building elements, or on panoramic images prone to severe distortions, misaligning with actual observations. Second, while various SVI-based building datasets have been established, their dependence on data derived from proprietary platforms introduces limitations related to accessibility, transparency, and adaptability. The ambiguous licensing terms of such datasets further constrain their utility for diverse research applications and compromise the integrity of work built upon them, thereby hindering inclusivity within the research community (Helbich et al., 2024). In a recent trend, crowdsourced SVI platforms have garnered attention in urban studies by producing diverse, publicly accessible imagery. Examples include annotating points of interest (Zarbakhsh and McArdle, 2023), image status (Hou et al., 2024), human perception (Yang et al., 2025), and road surface type (Kapp et al., 2025). Among these, Hou et al. (2024) curate a manually labeled dataset to assess 10 million crowdsourced SVIs from 688 cities, enriched with metadata such as platform, weather, and lighting conditions, while Kapp et al. (2025) utilize OSM tags and ChatGPT-4o to label and amplify underrepresented road surface classes, resulting in 9,122 labeled images. These initiatives illustrate the potential of crowdsourced data for broad, inclusive urban analyses.